# Hamiltonian simulation-based quantum-selected configuration interaction for large-scale electronic structure calculations with a quantum computer

### v1

@@ Abstract
Quantum-selected configuration interaction (QSCI) is one of the most promising approaches for quantum chemical calculations with the current pre-fault tolerant quantum computers. In the conventional QSCI, the Slater determinants used for the wave function expansion are sampled by iteratively performing approximate wave function preparation and subsequent measurement in the computational basis, and then the subspace Hamiltonian matrix is diagonalized on a classical computer. In this approach, the preparation of a high-quality approximate wave function is necessary to compute total energies accurately. In this work, we propose a Hamiltonian simulation-based QSCI (HSB-QSCI) to avoid this difficulty. In the HSB-QSCI, the Slater determinants are sampled from quantum states generated by the real-time evolution of approximate wave functions. We provide numerical simulations for the spin-singlet ground state and the first excited spin-triplet state of oligoacenes (benzene, naphthalene, and anthracene), phenylene-1,4-dinitrene, and hexa-1,2,3,4,5-pentaene molecules; these results reveal that the HSB-QSCI is applicable not only to molecules where the Hartree–Fock provides a good approximation of the ground state, but also to strongly correlated systems with multiconfigurational characteristics (i.e., the case where preparing a high-quality approximate wave function is hard). We have also numerically confirmed that the HSB-QSCI is robust to approximation errors of the Hamiltonian simulation, such as Trotter errors and the truncation errors of Hamiltonian term by maximum locality in the localized molecular orbital basis. Hardware demonstrations of the HSB-QSCI are also reported for the hexa-1,2,3,4,5-pentaene molecule using 20 qubits IBM superconducting device. The differences between the HSB-QSCI energy and the CAS-CI value are at most 0.15 kcal mol-1, achieving chemical precision.
@@ I Introduction
Among the diverse topics in the field of quantum computing, quantum chemical calculations of atoms and molecules have attracted attention as a promising application of quantum computers. A method to perform the full-configuration interaction (full-CI) calculation using the quantum phase estimation (QPE) algorithm was proposed in 2005 [1], and proof-of-principle experiments for the full-CI/STO-3G of the H2 molecule were reported in 2010 [2, 3]. The quantum circuit for QPE-based full-CI is usually very deep, so highly accurate QPE demonstrations were limited to small models with a few qubits [4, 5, 6, 7, 8]. Recently a QPE-type algorithm has been demonstrated in models with up to 33 qubits [9], however, it is still challenging to compute total energies of larger molecules with chemical precision (an error less than 1 kcal mol-1) on quantum computers available today.
@@
As an alternative approach to quantum chemical calculations using current quantum devices, the quantum-selected configuration interaction (QSCI) method has been proposed [14]. In the QSCI, quantum computers are used to sample Slater determinants that are important contributors to the ground state wave function. This can be done by running a quantum circuit to prepare an approximate wave function and measuring the quantum state in the computational basis. The subspace Hamiltonian matrix is then constructed from the Slater determinants selected via quantum computation and diagonalized on a classical computer. In this approach, the quantum computer is anticipated to efficiently sample a polynomial number of important Slater determinants from an exponentially large Hilbert space, and the nature of the quantum states used for sampling determines the accuracy of the calculation. At the stage of preparing an approximate wave function in QSCI, we may employ VQE or adiabatic state preparation method [14]. To advance the former technique, combining an adaptive strategy for ansatz construction (ADAPT [15]) to enhance the efficiency of VQE [16], and using a simpler cost function for VQE parameter optimization [17] have also been proposed. However, these approaches may still face challenges in variational optimizations when implemented on quantum hardware. It should also be noted that in the latter approach [17], the VQE wave function with optimized parameters does not necessarily correspond to the minimum with respect to the given Hamiltonian.
@@
Another strategy to prepare the quantum state for QSCI is to perform the coupled cluster singles and doubles (CCSD) calculation on a classical computer and embed the CCSD wave function using the local unitary cluster Jastrow (LUCJ) ansatz [18]. By using the LUCJ ansatz and introducing an error mitigation technique called as a self-consistent configuration recovery (SCCR), which is inspired by the structure of chemistry problems, Robeldo-Moreno and coworkers reported the QSCI calculations on the IBM superconducting quantum processor and supercomputer “Fugaku” for the nitrogen molecule (N2) and the iron–sulfur clusters [2Fe–2S] and [4Fe–4S] with up to 77 qubits [19]. This approach has also been applied to interaction energy [20] and excited state calculations [21, 22]. The combination of the QSCI with density matrix embedding theory has also been reported [23]. While this approach appears to be successful, it is not clear whether it is superior to the method of sampling important Slater determinants from the CCSD wave function on a classical computer.
@@
In this work, we report a Hamiltonian simulation-based quantum state preparation for QSCI (HSB-QSCI), in which important Slater determinants are sampled from the quantum states after the real-time evolutions of initial wave functions. Our proposed approach has several advantages over existing methods: (1) A simple approximate wave function, such as the HF wave function, can be used as the initial wave function, and no variational optimization in VQE is required. (2) Different samples can be obtained by changing the duration of the evolution time. (3) The accuracy of the Hamiltonian simulation does not need to be very high. As we will show later, the first-order Trotter decomposition with the time length of a single Trotter step is sufficient to collect important Slater determinants for the organic molecules under study. In addition, Hamiltonian term truncation based on operator locality in the localized molecular orbital basis can help to reduce the computational cost. (4) Feedback of the QSCI results into quantum computations is possible; that is, Slater determinants found to be important in the QSCI wave function can be added to the initial wave functions. (5) Although short-time evolution can be accurately simulated on a classical computer, long-time evolution poses significant challenges for classical simulation. Note that methods of sampling basis states for Hamiltonian matrix diagonalization based on Hamiltonian simulation have already been investigated in the framework of quantum Krylov subspace algorithms [24, 25]. Unlike the quantum Krylov method, which computes and on a quantum computer by sampling, the HSB-QSCI method requires only measurements in the computational (Pauli-Z) basis and is thus less computationally demanding. Here, , where , is the initial wave function, and is the Hamiltonian of the system. In the context of Monte Carlo simulation, a quantum dynamics Hamiltonian Monte Carlo (qdHMC) method is also proposed [26], in which the proposal step is performed on a quantum computer by sampling after time evolution.
@@
As a proof-of-concept demonstration of the HSB-QSCI method, we performed numerical simulations for the spin-singlet ground state and the first excited spin-triplet state of oligoacenes [benzene (12 qubits), naphthalene (20 qubits), and anthracene (28 qubits)], phenylene-1,4-dinitrene (20 qubits), and hexa-1,2,3,4,5-pentaene (20 qubits) in both planar and twisted geometries. Our numerical simulations showed that the HSB-QSCI can calculate the total energy with chemical precision by sampling the Slater determinants with a few steps of Hamiltonian simulations. We also report hardware demonstrations of the HSB-QSCI for the lowest spin-singlet and spin-triplet states of hexa-1,2,3,4,5-pentaene using ibm_torino processor, with the aid of matrix product operator (MPO)-based classical compression of the quantum circuit for time evolution. The HSB-QSCI energies calculated on ibm_torino processor agree with the CAS-CI values within 0.15 kcal mol-1 of error. For the calculation of spin triplet states, the accuracy of the HSB-QSCI energy was further improved by adding two important Slater determinants in the HSB-QSCI wave function to the initial wave function of the Hamiltonian simulation.
@@ II Theory
The initial wave function should have an overlap with the target electronic state. The HF wave function is a reasonable choice for in the electronic ground state calculations of typical closed-shell singlet molecules in their equilibrium geometries, but need not be a single Slater determinant. As we will discuss below, using an appropriately prepared multiconfigurational wave function as the initial wave function can reduce the number of sampled Slater determinants while keeping the QSCI energy almost unchanged. By defining the initial wave function as
@@
The effectiveness of sampling in the HSB-QSCI is mainly controlled by the evolution time length , the number of time steps , and the initial wave function . Note that from Eq. (6) we expect that must be set longer for systems with smaller energy gaps with the excited states. In the HSB-QSCI, we can reconstruct using the information of the QSCI wave function by constructing a multiconfigurational wave function consisting of several Slater determinants that have large contributions in the QSCI wave function in the previous step. Performing Hamiltonian simulations with different initial wave functions and merging the measurement results to perform the QSCI is another option. The availability of such feedback from the subspace Hamiltonian diagonalization part on a classical computer to the state preparation part on a quantum computer is one of the important features of the HSB-QSCI method.
Note that the Slater determinants are not always eigenfunctions of the electron spin operator (spin eigenfunctions). In fact, open-shell Slater determinants with spin- unpaired electron(s) are not spin eigenfunctions. Here we have assumed that , where and are the numbers of spin- and spin- electrons, respectively. In order to make the QSCI wave function being spin eigenfunctions, we introduced a symmetry completion step before constructing the Hamiltonian matrix. In the symmetry completion step, open-shell Slater determinants that are missing to span the spin eigenfunction are added. For example, if the bit string corresponding to the Slater determinant is sampled (2, , , and 0 stand for the corresponding molecular orbital is doubly occupied, singly occupied by a spin- electron, singly occupied by a spin- electron, and unoccupied, respectively) but the is not, then we add .
In case when the Hamiltonian simulations are carried out on a current noisy quantum hardware, the measurements of the quantum state give bit strings that correspond to the Slater determinants with different numbers of electrons. The SCCR proposed recently [19] can be adopted to recover the Slater determinants with a correct number of electrons. In the SCCR, if the bit string obtained from the measurement has the wrong number of electrons, bits are flipped to recover the desired number of electrons, and the probability of bit flipping is determined from the occupation number of the molecular orbital calculated from the QSCI wave function of the previous step. This process is iterated until convergence or the iteration step reaches the predefined maximum steps. The SCCR is performed before symmetry completion.
@@ III Computational conditions
To demonstrate the HSB-QSCI, we performed numerical simulations on a classical computer. The target systems are the spin-singlet ground state and the first excited spin-triplet state of oligoacenes (benzene, naphthalene, and anthracene), and phenylene-1,4-dinitrene as a representative system of the open-shell singlet ground state, and hexa-1,2,3,4,5-pentaene, as shown in Figure 2. Geometry optimizations of oligoacenes and hexa-1,2,3,4,5-pentaene in planar geometry were performed at the B3LYP/6-31G* level of theory. The twisted geometry of hexa-1,2,3,4,5-pentaene is generated by rotating one of the CH2 moieties 90 degrees from the equilibrium geometry. The geometry optimization of phenylene-1,4-dinitrene was carried out at the CASSCF(10e,10o)/6-31G* level. Here, (e,o) represents the active space consisting of electrons and molecular orbitals. Cartesian coordinates of all molecules are given in Supporting Information.
@@
Since the computational cost of numerical simulations of quantum circuits scales exponentially with the number of qubits, we adopted the active space approximation. We used the (6e,6o), (10e,10o), and (14e,14o) active spaces for benzene, naphthalene, and anthracene, which consist of the valence and orbitals. The one- and two-electron integrals of oligoacenes are calculated at the RHF/STO-3G level. The active space of phenylene-1,4-dinitrene contains the valence and orbitals and the in-plane 2p orbital of the nitrogen atoms, and the size is (10e,10o). The CASSCF(10e,10o)/6-31G* molecular orbitals are used to construct the second quantized Hamiltonian of phenylene-1,4-dinitrene. In the case of hexa-1,2,3,4,5-pentaene with (10e,10o) active space, the RHF/6-31G* calculation is performed and then the occupied molecular orbitals are localized using the Pipek–Mezey method [36]. For the virtual orbitals, we first formed the singular value decomposition (SVD) quasi-atomic external orbitals using an SVD with respect to the accurate atomic minimal basis functions, and then formed the ordered external orbitals using exchange integrals (using a keyword EXTLOC=ATMNOS in GAMESS-US software [37]). The active orbitals of all molecules are provided in Figures S1–S6 in Supporting Information. One- and two-electron atomic integrals are obtained from GAMESS-US, and the AO–MO integral transformations were performed using our own Python3 program. The qubit Hamiltonian is then generated using the OpenFermion library [38]. All the DFT calculations were done with the Gaussian 16 package [39], and the RHF and CASSCF calculations were carried out with the GAMESS-US software.
@@
As we mentioned in the previous section, the time evolution operation in the HSB-QSCI does not have to be exact, and we can introduce various approximations. In this work, we investigated the effect of Hamiltonian truncation on the sampled Slater determinants in the Hamiltonian simulation and on the QSCI energies. We investigated Hamiltonian truncation based on the locality of the qubit Hamiltonian terms. In this approach, we first generate the localized molecular orbitals (LMOs), and then rearrange the LMOs according to their relative spatial distances. During the qubit Hamiltonian construction using the reordered LMOs, we calculate the locality of the Pauli string (number of Pauli-X, Y, and Z operators in the Pauli string). If the locality of the Pauli string exceeds the threshold, then we exclude it from the qubit Hamiltonian. We investigated this method in hexa-1,2,3,4,5-pentaene, keeping in mind to perform the Hamiltonian simulations on a superconducting quantum device with MPO-based classical compression of the time evolution quantum circuit. The procedure of this step is illustrated in Figure 3.
@@
To sample the Slater determinants based on the real-time evolution of the initial wave function, we set the evolution time for to be and . To construct the quantum circuit for , we adopt the first-order Trotter decomposition with a single Trotter slice, unless otherwise specified. Magnitude ordering [40] is used for the Trotterized term ordering. The quantum circuit simulations were done using the qsim library [41], which allows us to use GPGPU. The number of shots for each is set to be unless otherwise noted. The QSCI part including SCCR, Hamiltonian matrix construction, and diagonalization was done using the qiskit-addon-sqd library [42].

### v2

@@ Abstract
Quantum-selected configuration interaction (QSCI) is an approach for quantum chemical calculations using current quantum computers. In conventional QSCI, Slater determinants used for the wave function expansion are sampled by iteratively performing approximate wave function preparation and subsequent measurement in the computational basis, and then the subspace Hamiltonian matrix is diagonalized on a classical computer. In this approach, preparation of a high-quality approximate wave function is necessary to accurately compute total energies. Here we propose a Hamiltonian simulation-based QSCI (HSB-QSCI) to avoid this difficulty, by sampling the Slater determinants from quantum states generated by the real-time evolution of approximate wave functions. We provide numerical simulations for the lowest spin-singlet and triplet states of oligoacenes (benzene, naphthalene, and anthracene), phenylene-1,4-dinitrene, and hexa-1,2,3,4,5-pentaene. We found that the HSB-QSCI is applicable not only to molecules where the Hartree–Fock provides a good approximation of the ground state, but also to strongly correlated systems where preparing a high-quality approximate wave function is hard. Hardware demonstrations of the HSB-QSCI are also reported for carbyne molecules expressed by up to 36 qubits, using an IBM Quantum processor. The HSB-QSCI captures more than 99.75% of the correlation energies in the active space by considering about 1% of all the Slater determinants in 36 qubit systems, illustrating the ability of the proposed method to efficiently consider important electronic configurations.
@@ I Introduction
Among the diverse topics in the field of quantum computing, quantum chemical calculations of atoms and molecules have attracted attention as a promising application of quantum computers. A method to perform the full-configuration interaction (full-CI) calculation using the quantum phase estimation (QPE) algorithm was proposed in 2005 [1], and proof-of-principle experiments for the full-CI/STO-3G of the H2 molecule were reported in 2010 [2, 3]. The quantum circuit for QPE-based full-CI is usually very deep, so highly accurate QPE demonstrations were limited to small models with a few qubits [4, 5, 6, 7, 8]. Recently a QPE-type algorithm has been demonstrated in models with up to 33 qubits [9], however, it is still challenging to compute total energies of larger molecules with chemical precision (an error less than 1.0 kcal mol-1) on quantum computers available today.
@@
As an alternative approach to quantum chemical calculations using current quantum devices, the quantum-selected configuration interaction (QSCI) method has been proposed [14]. In the QSCI, quantum computers are used to sample Slater determinants that are important contributors to the ground state wave function. This can be done by running a quantum circuit to prepare an approximate wave function and measuring the quantum state in the computational basis. The subspace Hamiltonian matrix is then constructed from the Slater determinants selected via quantum computation and diagonalized on a classical computer. In this approach, the quantum computer is anticipated to efficiently sample a polynomial number of important Slater determinants from an exponentially large Hilbert space, and the nature of the quantum states used for sampling determines the accuracy of the calculation. At the stage of preparing an approximate wave function in QSCI, we may employ VQE or the adiabatic state preparation method [14]. To advance the former technique, combining an adaptive strategy for ansatz construction (ADAPT [15]) to enhance the efficiency of VQE [16], and using a simpler cost function for VQE parameter optimization [17] have also been proposed. However, these approaches may still face challenges in variational optimizations when implemented on quantum hardware. It should also be noted that in the latter approach [17], the VQE wave function with optimized parameters does not necessarily correspond to the minimum with respect to the given Hamiltonian.
@@
Another strategy to prepare the quantum state for QSCI is to perform the coupled cluster singles and doubles (CCSD) calculation on a classical computer and embed the CCSD wave function using the local unitary cluster Jastrow (LUCJ) ansatz [18]. By using the LUCJ ansatz and introducing an error mitigation technique called as a self-consistent configuration recovery (SCCR), which is inspired by the structure of chemistry problems, Robledo-Moreno and coworkers reported the QSCI calculations on the IBM superconducting quantum processor and supercomputer “Fugaku” for the nitrogen molecule (N2) and the iron–sulfur clusters [2Fe–2S] and [4Fe–4S] with up to 77 qubits [19]. This approach has also been applied to interaction energy [20] and excited state calculations [21, 22]. The combination of the QSCI with density matrix embedding theory has also been reported [23]. While this approach appears to be successful, it is not clear whether it is superior to the method of sampling important Slater determinants from the CCSD wave function on a classical computer.
@@
In this work, we report a Hamiltonian simulation-based quantum state preparation for QSCI (HSB-QSCI), in which important Slater determinants are sampled from the quantum states after the real-time evolutions of initial wave functions 111After posting this manuscript to arXiv, Mikkelsen and Nakagawa, and Yu and coworkers independently submitted their own manuscripts to arXiv, presenting the same idea explored in this study. M. Mikkelsen, Y. O. Nakagawa, arXiv:2412.13839, 2024; J. Yu, et al, arXiv:2501.09721, 2025.. Our proposed approach has several advantages over existing methods: (1) A simple approximate wave function, such as the HF wave function, can be used as the initial wave function, and no variational optimization in VQE is required. (2) Different samples can be obtained by changing the duration of the evolution time. (3) The accuracy of the Hamiltonian simulation does not need to be very high. As we will show later, the first-order Trotter decomposition with the time length of a single Trotter step is sufficient to collect important Slater determinants for the organic molecules under study. In addition, Hamiltonian term truncation based on operator locality in the localized molecular orbital basis can help to reduce the computational cost. (4) Feedback of the QSCI results into quantum computations is possible; that is, Slater determinants found to be important in the QSCI wave function can be added to the initial wave functions. (5) Although short-time evolution can be accurately simulated on a classical computer, long-time evolution poses significant challenges for classical simulation.
Note that methods of sampling basis states for Hamiltonian matrix diagonalization based on Hamiltonian simulation have already been investigated in the framework of quantum Krylov subspace algorithms [25, 26]. Unlike the quantum Krylov method, which computes and on a quantum computer by sampling, the HSB-QSCI method requires only measurements in the computational (Pauli-Z) basis and is thus less computationally demanding. Here, , where , is the initial wave function, and is the Hamiltonian of the system. In the context of the Monte Carlo simulation, a quantum dynamics Hamiltonian Monte Carlo (qdHMC) method is also proposed [27], in which the proposal step is performed on a quantum computer by sampling after time evolution.
@@
As a proof-of-concept demonstration of the HSB-QSCI method, we performed numerical simulations for the spin-singlet ground (S0) state and the first excited spin-triplet (T1) state of oligoacenes [benzene (1; 12 qubits), naphthalene (2; 20 qubits), and anthracene (3; 28 qubits)], phenylene-1,4-dinitrene (4; 20 qubits), and hexa-1,2,3,4,5-pentaene (5; 20 qubits) in both planar and twisted geometries. Our numerical simulations showed that the HSB-QSCI can calculate the total energy with chemical precision by sampling the Slater determinants with about 10–20 steps of Hamiltonian simulations. We also report hardware demonstrations of the HSB-QSCI for the S0 and T1 states of three carbyne (one-dimensional sp-hybridized carbon atom chain) molecules including 5, octa-1,2,3,4,5,6,7-heptaene (6; 28 qubits), and deca-1,2,3,4,5,6,7,8,9-nonaene (7; 36 qubits) using an IBM Quantum processor ibm_kawasaki, with the aid of matrix product operator (MPO)-based classical compression of the quantum circuit for time evolution. The HSB-QSCI energies calculated on ibm_kawasaki agree with the CAS-CI values within 0.6 kcal mol-1 of error in 5. In 6, chemical precision for the total energy was not achieved with the real quantum device, but the HSB-QSCI predicted the singlet–triplet energy gap and the energy difference between planar and twisted geometries in chemical precision. The energy difference between planar and twisted geometries in the S0 state of 7 was also calculated with an error of 0.29 kcal mol-1. In 7, the HSB-QSCI captures more than 99.75% of correlation energies by considering only about 1% of all the Slater determinants.
@@ II Theory
The initial wave function should have an overlap with the target electronic state. The HF wave function is a reasonable choice for in the electronic ground state calculations of typical closed-shell singlet molecules in their equilibrium geometries, but need not be a single Slater determinant. By defining the initial wave function as
@@
The effectiveness of sampling in the HSB-QSCI is mainly controlled by the evolution time length , the number of time steps , and the initial wave function . Note that from Eq. (6), we expect that must be set longer for systems with smaller energy gaps with the excited states. In the HSB-QSCI, we can reconstruct using the information of the QSCI wave function by constructing a multiconfigurational wave function consisting of several Slater determinants that have large contributions in the QSCI wave function in the previous step. Performing Hamiltonian simulations with different initial wave functions and merging the measurement results to perform the QSCI is another option. The availability of such feedback from the subspace Hamiltonian diagonalization part on a classical computer to the state preparation part on a quantum computer is one of the important features of the HSB-QSCI method.
Note that the wave functions are simultaneous eigenfunctions of the Hamiltonian and electron spin operator in the non-relativistic regime, but Slater determinants are not always eigenfunctions of the operator (spin eigenfunctions). In fact, open-shell Slater determinants with spin- unpaired electron(s) are not spin eigenfunctions. Here we have assumed that , where and are the numbers of spin- and spin- electrons, respectively. In order to make the QSCI wave function being spin eigenfunctions, we introduced a symmetry completion step before constructing the subspace Hamiltonian matrix. In the symmetry completion step, open-shell Slater determinants that are missing to span the spin eigenfunction are added. For example, in the spin-singlet state calculations, if the bit string corresponding to the Slater determinant is sampled (2, , , and 0 mean that the corresponding molecular orbital is doubly occupied, singly occupied by a spin- electron, singly occupied by a spin- electron, and unoccupied, respectively), then we add , , , , and , if they are not sampled. Similarly for the spin-triplet state, if is measured, then we add , , and .
In a case when the Hamiltonian simulations are carried out on a current noisy quantum hardware, the measurements of the quantum state give bit strings that correspond to the Slater determinants with different numbers of electrons. The SCCR proposed recently [19] can be adopted to recover the Slater determinants with a correct number of electrons. In the SCCR, if the bit string obtained from the measurement has the wrong number of electrons, bits are flipped to recover the desired number of electrons, and the probability of bit flipping is determined from the occupation number of the molecular orbital calculated from the QSCI wave function of the previous step. This process is iterated until convergence or the iteration step reaches the predefined maximum steps. The SCCR is performed before symmetry completion.
@@ III Computational conditions
To demonstrate the HSB-QSCI, we performed numerical simulations on a classical computer. The target systems are the S0 and T1 states of oligoacenes (1, 2, and 3), 4 as a representative system of the open-shell singlet ground state, and 5, as shown in Figure 2. We also performed hardware demonstrations using an IBM Quantum processor for carbyne molecules 5, 6, and 7. Geometry optimizations of oligoacenes and carbynes in planar geometry were performed at the B3LYP/6-31G* level of theory. The twisted geometry of carbynes is generated by rotating one of the CH2 moieties 90 degrees from the equilibrium geometry. The geometry optimization of 4 was carried out at the CASSCF(10e,10o)/6-31G* level. Here, (e,o) represents the active space consisting of electrons and molecular orbitals. Cartesian coordinates of all molecules are given in Supporting Information.
@@
Since the computational cost of numerical simulations of quantum circuits scales exponentially with the number of qubits, we adopted the active space approximation. We used the (6e,6o), (10e,10o), and (14e,14o) active spaces for 1, 2, and 3, respectively, which consist of the valence and orbitals. The one- and two-electron integrals of oligoacenes are calculated at the RHF/STO-3G level. The active space of 4 contains the valence and orbitals and the in-plane 2p orbital of the nitrogen atoms, and the size is (10e,10o). The CASSCF(10e,10o)/6-31G* molecular orbitals are used to construct the second quantized Hamiltonian of 4. In the case of carbyne molecules 5, 6, and 7 with (10e,10o), (14e,14o), and (18e,18o) active spaces, respectively, the RHF/6-31G* calculation is performed and then the occupied molecular orbitals are localized using the Pipek–Mezey method [37]. For the virtual orbitals, we first formed the singular value decomposition (SVD) quasi-atomic external orbitals using an SVD with respect to the accurate atomic minimal basis functions, and then formed the ordered external orbitals using exchange integrals (using a keyword EXTLOC=ATMNOS in GAMESS-US software [38]). The active orbitals of all molecules are provided in Figures S1–S10 in Supporting Information. One- and two-electron atomic integrals are obtained from GAMESS-US, and the AO–MO integral transformations were performed using our own Python3 program. The qubit Hamiltonian is then generated using the OpenFermion library [39]. All the DFT calculations were done with the Gaussian 16 package [40], and the RHF and CASSCF calculations were carried out with the GAMESS-US software.
@@
As we mentioned in the previous section, the time evolution operation in the HSB-QSCI does not have to be exact, and we can introduce various approximations. In this work, we investigated the effect of Hamiltonian truncation on the sampled Slater determinants in the Hamiltonian simulation and on the QSCI energies. We investigated Hamiltonian truncation based on the locality of the qubit Hamiltonian terms. In this approach, we first generate the localized molecular orbitals (LMOs), and then rearrange the LMOs according to their relative spatial distances. During the qubit Hamiltonian construction using the reordered LMOs, we calculate the locality of the Pauli string (number of Pauli-X, Y, and Z operators in the Pauli string). If the locality of the Pauli string exceeds the threshold, then we exclude it from the qubit Hamiltonian. We investigated this method in 5, keeping in mind to perform the Hamiltonian simulations on a superconducting quantum device with MPO-based classical compression of the time evolution quantum circuit. The procedure of this step is illustrated in Figure 3.
@@
To sample the Slater determinants based on the real-time evolution of the initial wave function, we set the evolution time for to be and , unless otherwise noted. To construct the quantum circuit for , we adopt the first-order Trotter decomposition with a single Trotter slice. Magnitude ordering [41] is used for the Trotterized term ordering. The quantum circuit simulations were done using the qsim library [42], which allows us to use GPGPU. The subspace Hamiltonian matrix construction and diagonalization was done using PyCI library [43], and the SCCR for hardware demonstrations was carried out using qiskit-addon-sqd library [44]. It should be noted that the subspace Hamiltonian diagonalization can also be done using qiskit-addon-sqd, but it uses a selected CI kernel_fixed_space subroutine in PySCF [45] where the spin and occupancies are stored separately, and the Slater determinants constructed from all possible combinations of spin and strings are used as the basis for the wave function expansion. In this implementation, Slater determinants not sampled by the quantum computer can be included in the subspace Hamiltonian diagonalization, resulting in a quadratic increase in the dimension of the Hamiltonian matrix. This can be seen as a kind of subspace enlargement at the expense of compactness of the QSCI wave function. We also performed a subspace Hamiltonian diagonalization using PySCF, and the results are given in Supporting Information section D. Quantum circuit simulations and the subspace Hamiltonian diagonalization in the QSCI part were performed on the NVIDIA DGX H100 system.